OpenStudy (anonymous):

Does the Bernoulli effect work in en enclosed system (isolated system) where are Pascal's principles work, i.e. pressure transmits undiminished?

OpenStudy (anonymous):

I don't see why it shouldn't. Bernoulli's equation is applicable in a static fluid: imagine a set of lines running horizontal along a vertical pipe. Every point along each horizontal line has the same amount of energy (assuming a static liquid, \(v = 0\)), and therefore satisfies Bernoulli's equation, \[\frac{v^2}{2} + gh + \frac{p}{\rho} = \text{constant}\] along a streamline (a line of constant energy), while also fulfilling Pascal's law describing pressure between two points with a difference in height, h, \[ p_2 = p_1 + \rho g h \] I'm not quite sure if that answers your question but I hope it helps.